OpenStudy (anonymous):

9 - 4x + 7x = 15 A. x = -2 B. x = 8 C. x = -8

OpenStudy (anonymous):

this is easy so

OpenStudy (anonymous):

this is easy so

OpenStudy (anonymous):

Combine like terms: -4x + 7x = 3x 9 + 3x = 15

OpenStudy (anonymous):

Add '-9' to each side of the equation. 9 + -9 + 3x = 15 + -9

OpenStudy (anonymous):

First you combine like terms so -4x+7=3x. Then you get 9 +3x=15 then you subtract 9 from each side and then you will get 3x=6 then divide by 2. and it will b x=2

OpenStudy (anonymous):

Combine like terms: 9 + -9 = 0 0 + 3x = 15 + -9 3x = 15 + -9 Combine like terms: 15 + -9 = 6 3x = 6 Divide each side by '3'. x = 2 Simplifying x = 2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np